OpenStudy (anonymous):

Use a graphing utility to sketch the vectors and find the degree measure of the angle between the vectors 24. u = 2i - 3j v = 4i + 3j I did the graph but im not sure how to find the degree measure please help!

OpenStudy (callisto):

As I remember... \[cos\theta = \frac{x_1x_y+y_1y_2}{\sqrt{x_1^2+y_1^2}\sqrt{x_2^2+y_2^2}}\]So, \[cos\theta = \frac{2(4)+(-3)(3)}{\sqrt{2^2+(-3)^2}\sqrt{4^2+3^2}}\]Can you do it?

OpenStudy (anonymous):

yes! thank you!!!

OpenStudy (callisto):

Welcome :)